Wednesday, 24 June 2015

Maintenance Update (Henry, noticeboard, pony, hanger)

Weekend 20th and 21st
It was Thomas The Tank Engine weekend, so (as predicted) no one went to Toddington.  Anyway, 2807 was in service (still) as Henry.  Thomas was very poorly and had to be assisted by 4270 all of the time, so we were holding the fort between Cheltenham and Winchcombe all day each day.  On Sunday, Ray O'H (firing Thomas) said he was keeping him at 90 psi so that he could blow the whistle!  I noticed that his left-hand injector was permanently piddling.  The 04 shunter was playing with the Troublesome Trucks at Toddington, and laying a trail of sand up & down the track from its LH sandbox (it sure was, I was on it all day Saturday - Steve).  Driver SO threw a very public wobbly at a certain photographer who was standing in the middle of the track trying to get a full-on photo of Thomas approaching.  :-)

Gil & Fred went to Winchcombe on Saturday to continue with siphon restoration.  I took the time to re-do our noticeboard.

Wednesday 24th
Much to our surprise, 2807 was still in service!  There are a few new issues raised since last week:

32 [JCh]: LHS pony truck axle (inside) running hot. No oil in axlebox underkeep at end of day.

33 [JCh]: Vacuum pump release valves (both ends) hot.  {Now, we can't see how this can be, because they don't move (well, they open & close only) and the pump only pumps air.  So, they let air out to enable the creation of a partial vacuum.  So, "hot"???}

34 [JCh]: Pin holding reverser release handle to reverser has fallen out.  {replaced with a bolt.  Needs correct pin.}

35 [JCh]: Hole in front vacuum hose.  {Hose replaced}

36 [MY}: Rivets to both die blocks loose.  {We knew LHS one was loose, but only this winter replaced the RHS die block!}

37 [MY]: LH trailing tender underkeep T bolts loose.  {Nuts tightened & all checked}

38 [MY]: RH rear brake hanger (on loco) top pivot nuts loose.  {Nuts tightened.  Others OK}

As 2807 was still chuffing up & down, there was little to do but boot scraper production.  John G got stuck in to painting chairs - tops & bottoms.  By end of play, he'd painted about 14!

John T wielded the needle-gun and wire brush, cleaning random rail chairs.  Bruce assisted him during the afternoon.

I sanded and stained 12 brushes, and then applied black enamel to two rail chairs.

The two Johns and I assembled the noticeboard and fitted it back into position in the car park.  A gentleman from Milton Keynes was reading it as I left (so it was worth doing!).

Bruce investigated issues 27 & 32 concerning the pony truck during the morning, but despite clambering beneath the 42xx couldn't see what the cause of it running hot might be.  But until we can get beneath 2807 (when it stands still long enough!) it is hard to guess.

Oh, Gil was around (from time to time); but went for a ride on the footplate (ostensibly to check issues 32 & 33).  Also, Geof popped in to say "Hello", having discovered that there was no one working on the siphon today.

It would appear that 2807 is in service until the end of this month.  The 55xx will join it over this weekend (but was having a wash-out today).  So, the Saturday work party would best focus on the siphon restoration at Winchcombe … or else will be assigned to rail chair painting!


No comments: